OpenStudy (anonymous):

In a parallelogram, one angle is 9 times the size of another. Find the measures of the angles.

OpenStudy (anonymous):

LMNO is a parallelogram. If NM = x + 15 and OL = 3x + 5, find the value of x, then find NM and OL.

undeadknight26 (undeadknight26):

lol LMNO...laugh my ???? off?

undeadknight26 (undeadknight26):

any whoo i think the total its supposed to add up to is 180 so we have to find x to make it equal to 180...

OpenStudy (shrutipande9):

the sm of adjacent angles in parallelogram is 180....so take 1 as x and other as 9x

OpenStudy (anonymous):

what is answer

undeadknight26 (undeadknight26):

we can't just give you the answer sorry.

OpenStudy (shrutipande9):

exactly..work it out....9x+x=180

undeadknight26 (undeadknight26):

@shrutipande9 wouldn't it be 9x + x +15 or no?

OpenStudy (shrutipande9):

umm..no....i guess d comment is other part of the ques... first ques doesnt mention any 15

undeadknight26 (undeadknight26):

oh ok.

OpenStudy (shrutipande9):

 Opposite sides of a parallelogram are congruent. OL and NM are opposite sides. so x+15=3x+5

OpenStudy (shrutipande9):

i hope u can solve it further

OpenStudy (anonymous):

i get it

OpenStudy (anonymous):

Find the values of x and y for which ABCD must be a parallelogram.